OpenStudy (anonymous):

factor x^2 – 8x + 15

OpenStudy (ghazi):

x^2-5x-3x+15 =(x-5)(x-3)

OpenStudy (anonymous):

can you show the steps inbetween? I want to understand how to do it

OpenStudy (ghazi):

okay , just break -8x in such a way so that the product of those two numbers should be equal to 15

OpenStudy (ghazi):

when i multiply -5 to -3 , i will get +15

OpenStudy (ghazi):

and also sum of -5 and -3 = -8

OpenStudy (ghazi):

therefore write it like \[x^2-5x-3x+15=x(x-5)-3(x-5)=(x-3)(x-5)\] keep taking out common values